the sum of 2 numbers is 74.when the bigger number is divided by the smaller number,the answer is 5 remainder 8. what are the two numbers?
Answer by privani(34)

Assuming the bigger number is x, and the other number is y.
Form equations from Q:
x+y=74 ... (1)
x/y=5 remainder 8
x=5y+8 ... (2)
Substitute (2) in (1):
x+y=74
5y+8+y=74
6y+8=74
6y=74-8
y=66/6
y=11
Substitute y=11 in (1):
x+y=74
x+11=74
x=74-11
x=63